OpenStudy (anonymous):

what is the sum of the geometric sequence -3, 21, 147 if there are 6 terms

OpenStudy (amistre64):

since it is a geometric sequence, there is some common ratio such that\[a_n=a_{n-1}~r\]

OpenStudy (amistre64):

are you by chance missing a negative on that 147?

OpenStudy (anonymous):

@amistre64 yep! sorry :(

OpenStudy (amistre64):

\[a_n=a_{n-1}~r\] in otherwords; our common ratio is such that -3 * r = 21 we need this to find the sum of all the terms required

OpenStudy (amistre64):

\[S_n=a_1\frac{1-r^n}{1-r}\] \[S_6=-3\frac{1-r^6}{1-r}\]